# ATLAS: Automated Tree-based Language Analysis System for C and C++ Source Programs

### v2

@@ Abstract.
Analyzing non-compilable C/C++ submodules without a resolved build environment remains a critical bottleneck for industrial software evolution. Traditional static analysis tools often fail in these scenarios due to their reliance on successful compilation, while Large Language Models (LLMs) lack the structural context necessary to reason about complex program logic. We introduce ATLAS, a Python-based Command-Line Interface (CLI). It is characterized by: (i) inter-procedural, type-aware analysis that captures dependencies across function boundaries; (ii) support for both full-project and partial analysis of non-compilable projects; (iii) graph optimizations, such as variable collapsing and node blacklisting; and (iv) synchronized multi-view graphs that align syntax, execution paths, and data-flow logic. Evaluating ATLAS with DeepSeek V3.2 for automated test generation demonstrates a 34.71% increase in line coverage and 32.66% in branch coverage, achieving performance on par with and even exceeding the industry-standard symbolic execution tool KLEE on complex projects. With polynomial scalability, ATLAS provides a robust infrastructure for generating the information-dense datasets required by next-generation, graph-aware Machine Learning for Software Engineering (ML4SE) models.
Video demonstration: https://youtu.be/xEJDNYkiD4s
Tool github repository: https://github.com/jaid-monwar/ATLAS-multi-view-code-representation-tool.git
@@ 1. Introduction
Programming languages like C and C++ are considered still as the foundation of performance-critical software, such as operating systems and embedded firmware (Stroustrup, 2020), where hardware proximity and resource efficiency are paramount. However, as C/C++ based systems scale, managing their evolution becomes increasingly complex (Brooks, 1987), particularly when analyzing sub-modules that are non-compilable or contain unresolved dependencies. Consequently, the field has transitioned toward the Machine Learning for Software Engineering (ML4SE) to automate tasks like test generation (Fontes and Gay, 2023) (Durelli et al., 2019), code translation (Tufano et al., 2019) (Ibrahimzada et al., 2025), source code classification (Mou et al., 2016), and code summarization (Haiduc et al., 2010) (Jiang et al., 2017). Within this paradigm, Large Language Models (LLMs) offer a flexible alternative to traditional tools. These can derive semantic context even from unbuildable projects by processing code as structured textual sequences (Wang et al., 2025) (Panichella et al., 2013) (Li et al., 2025).
However, LLMs are limited by their probabilistic nature, often prioritizing token adjacency over the deep, non-linear relationships that dictate program behavior (Anand et al., 2024) (Pendyala and Thakur, 2025). Bridging this ’semantic gap’ requires effective representation learning. Such techniques are essential for training and evaluating ML models while also enhancing the reasoning capabilities of LLMs. Empirical evidence (Liu and Jabbarvand, 2025) (Samoaa et al., 2022) shows that performance improves significantly when models utilize multi-view data that explicitly encodes these underlying logical relationships.
To bridge this gap, research has shifted toward representing code as integrated graphs (Chimalakonda et al., 2023). The Abstract Syntax Tree (AST) encodes the hierarchical syntactic structure of code constructs, the Control Flow Graph (CFG) models execution paths, and the Data Flow Graph (DFG) captures the variable dependencies. However, their integration into a multi-view representation allows models to holistically capture program logic. This approach is validated by CodeSAM (Mathai et al., 2024), which uses attention-masking to improve code search, and PANTA (Gu et al., 2025), which leverages CFG analysis to increase the test coverage by 23%. However, these successes are largely centered on Java, and C/C++ projects lack such high-level abstractions and require significantly more code to implement equivalent logic, exhausting LLM context windows more rapidly(Chimalakonda et al., 2023). Moreover, this difficulty is further complicated by C-specific hurdles like pointer aliasing, and type obscuration, which remain under-addressed (Schubert et al., 2021) in representation learning research.
@@ 2. Related Work
Additionally, static tools like Joern (14) provide multi-language code-views but lack a native mechanism to customize or combine these representations into a unified structure. This often results in isolated, function-level graphs that cannot resolve the long-range dependencies required for a comprehensive project-level analysis. Similarly, while Tree-sitter (Brunsfeld and Contributors, 2023) parses syntax, it cannot resolve execution logic or data-flow. In contrast, ATLAS employs custom inter-procedural algorithms to precisely model pointer indirection and class-member state at the statement level. While COMEX (Das et al., 2023) targets memory-safe Java and C#, ATLAS addresses unmanaged C/C++ semantics through specialized, type-aware analysis.
ATLAS further addresses two critical gaps in the current landscape. First, it provides statement-level DFGs for C and C++, whereas existing research tools like PhASAR (Schubert et al., 2019) focus primarily on IR. Second, it resolves the fragmentation of current toolchains, where developers must currently stitch together disparate tools, such as Clang for ASTs and Joern (14) for CFGs, using custom scripts. This fragmented process is inefficient and lacks cross-view alignment. ATLAS provides a unified framework for aligned multi-view graphs, eliminating redundant analysis and streamlining ML4SE dataset pipelines.
@@ 3. ATLAS Tool
ATLAS follows a two-step workflow (Fig. 1). At first, the Preprocessor and Parser normalize multi-file projects into a Parsed Code Model enriched with scope, type, and declaration-use metadata. Second, the Combined Code View Generator constructs aligned graphs via four modules: the AST, CFG, and DFG Drivers, and the Unified Multi-view Graph. These components utilize Reaching Definition Analysis (RDA) and inter-procedural mapping to ensure consistent node alignment across all views. Finally, ATLAS exports these representations in JSON, DOT, or PNG, linking a program’s syntactic structure directly to its execution behavior.
@@ 3.1. Preprocessor and Parser
The Preprocessor and Parser generate a unified Parsed Code Model. The Preprocessor normalizes the codebase by consolidating multi-file projects, removing non-semantic elements (e.g., comments and empty lines), and resolving global macro and type definitions. Subsequently, the Parser leverages Tree-sitter to generate concrete syntax trees (CSTs), which are traversed to build an enriched symbol table mapping scope hierarchies, data types, and declaration-use relationships. This metadata layer serves as the foundational reference for all subsequent graph generation.
@@ 3.2. Code Representation Generation
The Combined Code View Generator utilizes the enriched symbol table to construct aligned AST, CFG, and DFG views. This shared reference ensures node-level consistency for unified multi-view analysis. Construction logic and refinements for each representation are as follows.
@@ 3.2.1. AST
The CST produced in the preprocessing step preserves all syntactic details, which ATLAS traverses to build an AST index mapping each node’s position and type to a unique identifier. To refine this representation, ATLAS prunes non-semantic tokens, such as brackets and semicolons, to produce a core AST, as shown in Fig. 2(a). The AST Driver further provides two customization mechanisms to control graph granularity: (i) Variable Collapsing, which unifies all occurrences of a specific variable identifier into a single node to reduce redundancy, and (ii) Node Blacklisting, which allows users to omit irrelevant node categories (e.g., comment, modifier) and their children. These refinements allow users to distill the program into its core logical elements, effectively mitigating the computational overhead and overfitting risks associated with high-dimensional syntax.
@@ 3.2.2. CFG
The CFG Driver generates statement-level, inter-procedural representations where each node corresponds to a source statement and edges map all potential execution transitions. Unlike traditional tools that produce isolated function-level graphs, ATLAS establishes global connectivity by resolving function calls as directed edges from caller statements to callee definitions. This capability is demonstrated in Fig. 2(b), where the ternary operator on line 20 triggers indirect calls to f1() and f2() via function pointers; ATLAS correctly models this ambiguity by generating branching edges to both function definitions. Furthermore, the tool natively supports recursion by representing it as self-referential edges, as seen in the recursive call within f1() on line 7. By unifying these complex control structures into a single connected graph, ATLAS enables a comprehensive trace of program execution often lost in intra-procedural analysis.
@@ 3.2.4. Combined Multi-View Graph
ATLAS merges these complementary views into a unified multi-view graph by preserving consistent node identities across all representations. This alignment allows downstream ML/DL models to learn from syntactic, control-flow, and data-flow properties simultaneously. The tool supports fifteen combination variants, including pairwise and triple-view integrations, all of which remain compatible with user-defined collapsing and blacklisting. By assigning distinct type attributes to each edge class, ATLAS enables models to selectively prioritize specific program semantics over detail-heavy syntax.
@@ 3.3. Visualizer
The Visualizer manages the exportation of aligned graphs into three different formats. While the JSON output is optimized for Machine Learning (ML) integration, the DOT and PNG formats facilitate human-centric debugging and structural inspection. For multi-view representations, the component employs color-coded edge classes to distinguish between syntactic, control, and data-flow properties. Additionally, the Visualizer implements specialized C++ formatting, such as escaping namespace qualifiers and reserved keywords, to ensure high-fidelity rendering of complex codebase structures.
@@ 4. ATLAS In Practice
Table 2. ATLAS Supported Language Constructs
Functions
Loops/ Conditionals
Struct/Enum/ Union
Func. Pointers
Goto/Label
Classes
Namespaces
Templates
Virtual Methods
Exceptions
Op. Overloading
C
C++
@@
To evaluate the accuracy of ATLAS, we generate and verify manually the CFG and DFG for all the C and C++ projects from the TheAlgorithms dataset (TheAlgorithms, 2025b) (TheAlgorithms, 2025a). To evaluate the practicality of the tool, we compare the test generation performance of DeepSeek V3.2 with and without statement-level CFG as context. Table 2 summarizes the language constructs natively supported by ATLAS. We select the 10 largest C programs by LOC, excluding main functions and test scaffolding to isolate the core algorithmic logic. Additionally, we benchmark the tool’s scalability by monitoring its computational resource consumption relative to code complexity.
4.1. Impact of CodeView Combinations
Our analysis lies in advancing LLM-based control-flow reasoning for test generation, therefore our focus is on the comparison between state-of-the-art LLM with statement-level CFG (DS w/ CFG) and without CFG (DS w/o CFG). We use DeepSeek V3.2 with temperature=0, max_tokens=4096, and reasoning turned off. For DS w/ CFG, we extract the execution paths from the statement-level CFG and generate a test for each unique path. We replicate a similar architecture to PANTA with max_iterations=15 and LLM temperature=0.1 for test repairs. We discard any tests if they still fail after max iterations. For DS w/o CFG, we do not add the paths into the LLM’s context as we do not generate a CFG, and we keep the rest of the modules exactly identical.
We present the results in Table 1. In terms of line and branch coverages, DS w/ CFG consistently outperforms DS w/o CFG across all projects. On average, DS w/ CFG achieves 34.71% higher line coverage and 32.66% higher branch coverage than DS w/o CFG. Moreover, the performance is on-par with the industry-standard tool KLEE (Cadar et al., 2008) in most projects and even exceeds it in the larger projects: red_black_tree, multikey_quick_sort, hangman, and trie. Beyond coverage metrics, DS w/ CFG also demonstrates improved test efficiency: 75% of generated tests exercise unique execution paths compared to only 67% for DS w/o CFG. This indicates that CFG-guided test generation reduces redundant tests by enabling the LLM to target distinct control-flow paths systematically.
4.2. Scalability Analysis
To evaluate the scalability of ATLAS, we measure its performance during CFG generation on the 20 largest programs from the TheAlgorithms C dataset (TheAlgorithms, 2025b), ranging from 61 to 309 LOC. Figure 3 presents the execution time and memory consumption metrics across these programs. The median execution time is 6.89 seconds, with median peak memory usage of 135.01 MB. The trend lines in Figure 3 show best-fit polynomial curves, revealing that execution time and memory consumption scale polynomially with program size rather than exponentially, suggesting that ATLAS remains practical for moderately-sized programs typical of benchmark datasets.
5. Discussion and Limitations
ATLAS provides robust C/C++ support by prioritizing core program logic and inter-procedural dependencies. By requiring only syntactic validity rather than full compilability, the tool bypasses build-system overhead while maintaining broad industrial applicability. To preserve source-level context, ATLAS parses pre-processor directives without executing conditional compilation paths, ensuring it remains a flexible analyzer rather than a rigid compiler replacement. Furthermore, certain advanced C++ features, such as template metaprogramming, SFINAE, and complex inheritance hierarchies, are currently modeled with limited granularity to maintain graph clarity and computational scalability. These design choices ensure that ATLAS remains a performant infrastructure for generating the information-dense datasets required by modern ML4SE models.
6. Conclusion
We introduced ATLAS, a CLI that generates unified multi-view representations of C/C++ projects by aligning syntactic, control, and data-flow properties. By operating on raw source code without build-system dependencies, ATLAS enables the rapid generation of information-dense datasets for ML4SE. Our evaluation proves that ATLAS-guided LLMs significantly enhance control-flow reasoning, matching the effectiveness of symbolic execution tools like KLEE with polynomial scalability. ATLAS establishes a foundational infrastructure for graph-aware models, with future work targeting advanced C++ idioms and automated memory-safety analysis.
Appendix A ATLAS: Tool Demonstration Walkthrough
This appendix presents an evaluation of ATLAS’s success rates, compares its CFG representation with LLVM IR, and provides step-by-step demonstrations showcasing single-file analysis, combined multi-code view generation, and multi-file project analysis.
A.1. CFG and DFG Generation Success Rate
@@
We evaluated ATLAS on the C and C++ subsets of The Algorithms datasets. The C dataset contains 406 files, where ATLAS achieved a success rate of 96.80% for CFG and 91.38% for DFG generation. The C++ dataset consists of 360 files, with success rates of 91.67% for CFGs and 90.56% for DFGs. The remaining uncovered cases are due to specific, complex language features that are not yet fully supported. Table 3 provides a detailed breakdown of the failure categories for both languages and analysis types.
A.2. CFG Representation Comparison: LLVM IR vs. ATLAS

### v3

@@ Abstract.
Analyzing non-compilable C/C++ submodules without a resolved build environment remains a critical bottleneck for industrial software evolution. Traditional static analysis tools often fail in these scenarios due to their reliance on successful compilation, while Large Language Models (LLMs) lack the structural context necessary to reason about complex program logic. We introduce ATLAS, a Python-based CLI that generates unified multi-view representations for large-scale C/C++ projects with high accuracy, achieving success rates up to 96.80% for CFGs and 91.38% for DFGs. ATLAS is characterized by: (i) inter-procedural, type-aware analysis across function boundaries; (ii) support for both full and partial analysis of non-compilable projects; (iii) graph optimizations such as variable collapsing and node blacklisting; and (iv) synchronized multi-view graphs that align syntax, execution paths, and data-flow logic. Evaluating ATLAS with DeepSeek V3.2 for automated test generation demonstrates a 34.71% increase in line coverage and 32.66% in branch coverage, matching or exceeding the performance of the symbolic execution tool KLEE on complex projects. With polynomial scalability, ATLAS provides a robust infrastructure for generating the information-dense datasets required by next-generation, graph-aware ML4SE models.
Video demonstration: https://youtu.be/QGuJZhj9CTA
Tool Github repository: ATLAS
@@ 1. Introduction
Programming languages like C and C++ are still considered the foundation of performance-critical software, such as operating systems and embedded firmware (Stroustrup, 2020), where hardware proximity and resource efficiency are paramount. However, as C/C++ based systems scale, managing their evolution becomes increasingly complex (Brooks, 1987), particularly when analyzing sub-modules that contain unresolved dependencies. Consequently, the field has transitioned towards Machine Learning for Software Engineering (ML4SE) to automate tasks like test generation (Fontes and Gay, 2023) (Durelli et al., 2019), code translation (Tufano et al., 2019) (Ibrahimzada et al., 2025), source code classification (Mou et al., 2016), and code summarization (Haiduc et al., 2010) (Jiang et al., 2017). Within this paradigm, Large Language Models (LLMs) offer a flexible alternative to traditional tools. LLMs can derive semantic context even from unbuildable projects by processing code as structured textual sequences (Wang et al., 2025) (Panichella et al., 2013) (Li et al., 2025).
However, LLMs are limited by their probabilistic nature, often prioritizing token adjacency over the deep, non-linear relationships that dictate program behavior (Anand et al., 2024) (Pendyala and Thakur, 2025). Bridging this ’semantic gap’ requires effective representation learning. Such techniques are essential for training and evaluating Machine Learning (ML) models while also enhancing the reasoning capabilities of LLMs. Empirical evidence (Liu and Jabbarvand, 2025) (Samoaa et al., 2022) shows that performance improves significantly when models utilize multi-view data that explicitly encodes these underlying logical relationships.
Consequently, research has shifted toward representing code as integrated graphs (Chimalakonda et al., 2023). The Abstract Syntax Tree (AST) encodes the hierarchical syntactic structure of code constructs, the Control Flow Graph (CFG) models execution paths, and the Data Flow Graph (DFG) captures the variable dependencies. Crucially, their integration into a multi-view representation allows models to holistically capture program logic. This approach is validated by CodeSAM (Mathai et al., 2024), which uses attention-masking to improve code search, and PANTA (Gu et al., 2025), which leverages CFG analysis to increase the test coverage by 23%. However, these successes are largely centered on Java. C/C++ projects, on the other hand, lack such high-level abstractions and require significantly more code to implement equivalent logic, exhausting LLM context windows more rapidly (Chimalakonda et al., 2023). Moreover, this difficulty is further complicated by C-specific hurdles like pointer aliasing and type obscuration, which remain under-addressed (Schubert et al., 2021) in representation learning research.
@@ 2. Related Work
Additionally, static tools like Joern (joe, [n. d.]) provide multi-language code-views but lack a native mechanism to customize or combine these representations into a unified structure. This often results in isolated, function-level graphs that cannot resolve the long-range dependencies required for a comprehensive project-level analysis. Similarly, while Tree-sitter (Brunsfeld and Contributors, 2023) parses syntax, it cannot resolve execution logic or data-flow. In contrast, ATLAS employs custom inter-procedural algorithms to precisely model pointer indirection and class-member state at the statement level. While COMEX (Das et al., 2023) targets memory-safe Java and C#, ATLAS addresses unmanaged C/C++ semantics through specialized, type-aware analysis.
ATLAS further addresses two critical gaps in the current landscape. First, it provides statement-level DFGs for C and C++, whereas existing research tools like PhASAR (Schubert et al., 2019) focus primarily on IR. Second, it resolves the fragmentation of current toolchains, where developers must stitch together disparate tools, such as Clang for ASTs and Joern (joe, [n. d.]) for CFGs, using custom scripts. This fragmented process is inefficient and lacks cross-view alignment. ATLAS provides a unified framework for aligned multi-view graphs, eliminating redundant analysis and streamlining ML4SE tasks.
@@ 3. ATLAS Tool
ATLAS follows a three-step workflow (Fig. 1). First, the Preprocessor and Parser normalize multi-file projects into a Parsed Code Model enriched with scope, type, and declaration-use metadata. Second, the Combined Code View Generator constructs aligned graphs via four modules: the AST, CFG, and DFG Drivers, and the Unified Multi-view Graph. Finally, ATLAS exports these representations in JSON, DOT, or PNG, linking a program’s syntactic structure directly to its execution behavior.
@@ 3.1. Preprocessor and Parser
The module generates a unified Parsed Code Model. The Preprocessor normalizes the codebase by consolidating multi-file projects, removing non-semantic elements (e.g., comments and empty lines), and resolving global macro and type definitions. Subsequently, the Parser leverages Tree-sitter to generate concrete syntax trees (CSTs), which are traversed to build an enriched symbol table mapping scope hierarchies, data types, and declaration-use relationships.
@@ 3.2. Code Representation Generation
The Combined Code View Generator utilizes the enriched symbol table to construct aligned AST, CFG, and DFG views. These components utilize Reaching Definition Analysis (RDA) and inter-procedural mapping to ensure consistent node alignment across all views. Construction logic and refinements for each representation are as follows.
@@ 3.2.1. AST
The CST produced in the preprocessing step preserves all syntactic details, which ATLAS traverses to build an AST index, mapping each node’s position and type to a unique identifier. To refine this representation, ATLAS prunes non-semantic tokens, such as brackets and semicolons, to produce a core AST (Fig. 2(a)). The AST Driver further provides two customization mechanisms to control graph granularity: (i) Variable Collapsing, which unifies all occurrences of a specific variable identifier into a single node to reduce redundancy, and (ii) Node Blacklisting, which allows users to omit irrelevant node categories (e.g., modifier) and their children. These refinements allow users to distill the program into its core logical elements, effectively mitigating the computational overhead and overfitting risks associated with high-dimensional syntax.
@@ 3.2.2. CFG
The CFG Driver generates statement-level and inter-procedural representations where each node corresponds to a source statement and edges map all potential execution transitions. Unlike traditional tools that produce isolated function-level graphs, ATLAS establishes global connectivity by resolving function calls as directed edges from caller statements to callee definitions. This capability is demonstrated in Fig. 2(b), where the ternary operator on line 20 triggers indirect calls to f1() and f2() via function pointers. ATLAS correctly models this ambiguity by generating branching edges to both function definitions. Furthermore, the tool natively supports recursion by representing it as self-referential edges, as seen in the recursive call within f1() on line 7. By unifying these complex control structures into a single connected graph, ATLAS enables a comprehensive trace of program execution often lost in intra-procedural analysis.
@@ 3.2.4. Combined Multi-View Graph
ATLAS merges AST, CFG, and DFG views into a unified multi-view graph by preserving consistent node identities across all representations. This alignment enables downstream models to learn syntactic, control-flow, and data-flow properties simultaneously. The tool supports fifteen combination variants, including pairwise and triple-view integrations, fully compatible with user-defined collapsing and blacklisting. By assigning distinct type attributes to each edge class, ATLAS allows models to prioritize specific program semantics over detail-heavy syntax.
@@ 3.3. Visualizer
The Visualizer manages the exportation of aligned graphs into three different formats. While the JSON output is optimized for ML integration, the DOT and PNG formats facilitate human-centric debugging and structural inspection. For multi-view representations, the component employs color-coded edge classes to distinguish between syntactic, control, and data-flow properties.
@@ 4. ATLAS In Practice
To evaluate the accuracy of ATLAS, we generate and manually verify the CFG and DFG for all the C and C++ projects from the TheAlgorithms dataset (TheAlgorithms, 2025) (TheAlgorithms, 2025). To evaluate the practicality of the tool, we compare the test generation performance of DeepSeek V3.2 (DeepSeek-AI et al., 2025) with and without statement-level CFG as context. We select the 10 largest C programs by LOC, excluding main functions and test scaffolding to isolate the core algorithmic logic. Additionally, we benchmark the tool’s scalability by monitoring its computational resource consumption relative to code complexity.
4.1. Graph Generation Success Rates
@@
Evaluating ATLAS on the TheAlgorithms dataset (406 C and 360 C++ files) yielded high success rates for graph generation: CFG: 96.80% (C) and 91.67% (C++); DFG: 91.38% (C) and 90.56% (C++). As detailed in Table 2, remaining failures stem from the unmanaged memory model of C/C++, where runtime behaviors, such as dynamic pointer arithmetic and array element modifications, cannot be fully resolved statically. Additionally, to prioritize graph clarity and computational scalability, certain complex constructs like goto statements and static class variables are currently modeled with limited granularity. By requiring only syntactic validity rather than a resolved build environment, ATLAS bypasses the overhead of compilation while maintaining broad industrial applicability. These design trade-offs establish ATLAS as a performant tool for generating the information-dense datasets essential for ML models.
4.2. Impact of CodeView Combinations
Our analysis lies in advancing LLM-based control-flow reasoning for test generation; therefore, our focus is on the comparison between state-of-the-art LLM with statement-level CFG (DS w/ CFG) and without CFG (DS w/o CFG). We use DeepSeek V3.2 with temperature=0, max_tokens=4096, and reasoning turned off. For DS w/ CFG, we extract the execution paths from the statement-level CFG and generate a test for each unique path. We replicate a similar architecture to PANTA (Gu et al., 2025) with max_iterations=15 and LLM temperature=0.1 for test repairs. We discard any tests if they still fail after max iterations. For DS w/o CFG, we do not add the paths into the LLM’s context as we do not generate a CFG, and we keep the rest of the modules exactly identical.
We present the results in Table 1. In terms of line and branch coverages, DS w/ CFG consistently outperforms DS w/o CFG almost across all projects. On average, DS w/ CFG achieves 34.71% higher line coverage and 32.66% higher branch coverage than DS w/o CFG. The coverage performance of DS w/ CFG is on-par with the industry-standard tool KLEE (Cadar et al., 2008) in most projects and even exceeds it in the larger projects (red_black_tree, hangman, etc). Beyond coverage metrics, DS w/ CFG also demonstrates improved test efficiency of around 75% in terms of the ratio of tests that follow at least one unique execution path to the total number of generated tests. This is in stark contrast to only 67% for DS w/o CFG. This indicates that ATLAS-generated statement-level CFG-guided test generation reduces redundant tests by enabling the LLM to target distinct control-flow paths systematically.
4.3. Scalability Analysis
To evaluate the scalability of ATLAS, we measure its performance during CFG generation on the 20 largest C programs from the TheAlgorithms C dataset (TheAlgorithms, 2025), ranging from 61 to 309 LOC. Figure 3 presents the execution time and memory consumption metrics across these programs. The median execution time is 6.89 seconds, with median peak memory usage of 135.01 MB. The trend lines in Figure 3 show best-fit polynomial curves, revealing that execution time and memory consumption scale polynomially with program size rather than exponentially, suggesting that ATLAS remains practical for moderately-sized programs typical of benchmark datasets.
5. Conclusion
We introduced ATLAS, a Python-based CLI that generates unified multi-view representations of C/C++ projects by aligning syntactic, control, and data-flow properties. By operating on raw source code without build-system dependencies, ATLAS enables the rapid generation of information-dense datasets for ML4SE. Our evaluation demonstrates that ATLAS-guided LLMs significantly enhance control-flow reasoning, matching the effectiveness of symbolic execution tools like KLEE, while maintaining polynomial scalability. ATLAS establishes a foundational infrastructure for graph-aware models, with future work targeting advanced C++ idioms and automated memory-safety analysis.
Appendix A ATLAS: Tool Demonstration Walkthrough
This appendix compares ATLAS’s CFG representation with LLVM IR and provides step-by-step demonstrations showcasing single-file analysis, combined multi-code view generation, and multi-file project analysis.
Video demonstration: https://youtu.be/QGuJZhj9CTA
Tool Github repository: ATLAS
A.1. CFG Representation Comparison: LLVM IR vs. ATLAS